# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ Appendix A Proofs in Section 2
$p_{T}({\omega}^{\prime}\mid{\omega})=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},e^{{{\mathbf{G}}}T}{\hat{\bm{e}}}_{{\omega}}\rangle=\int{\mathcal{D}}{\omega}\ \exp\left\{-\sum_{x}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t,x},w_{t,x}^{+})\right\}.$ (A.70)
@@
It corresponds to an operator ${\bm{\Gamma}}:{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$ defined as
${\bm{\Gamma}}:=\gamma_{{\omega}^{\prime}{\omega}}{{\check{\bm{\eta}}}}^{{\omega}^{\prime}{\omega}},\qquad\text{where }\ \forall{\omega}:\gamma_{{\omega}{\omega}}=0.$ (A.72)
@@
$\mathbb{E}\left[\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \gamma(t,{\omega},{\omega}^{+})\Bigm{|}{\omega}_{0}={\omega}\right]=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}{\hat{\bm{e}}}_{{\omega}}.$ (A.73)
@@
$=\left(\gamma_{{\omega}^{\prime\prime}{\omega}}{{\check{\bm{\eta}}}}^{{\omega}^{\prime\prime}{\omega}}\right)\left({\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}}G^{{\omega}^{\prime}}_{{\omega}}\right)\cdot{\Delta t}+o({\Delta t})$ (A.76)
$={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\cdot{\Delta t}+o({\Delta t}).$ (A.77)
@@
$=\sum_{n=1}^{N}\sum_{{\omega}^{\prime}}\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},e^{(n-1){{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}}\rangle\cdot\left({\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}^{\prime}}\cdot{\Delta t}+o({\Delta t})\right)$ (A.80)
$=\sum_{n=1}^{N}{\Delta t}\cdot\left({\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{(n-1){{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}}+o({\Delta t})/{\Delta t}\right)$ (A.81)
$=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}{\hat{\bm{e}}}_{{\omega}}.$ (as ${\Delta t}\to 0$ and $N\to\infty$)
@@
$\tilde{\bm{L}}{\widetilde{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}=-K\log K+H_{\omega}.$ (A.83)
@@
For any ${\hat{\bm{e}}}_{{\omega}}\in{\mathcal{H}}$, we have
$\tilde{\bm{L}}{\widetilde{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}$ $=-\sum_{{\omega}^{\prime}:{\omega}^{\prime}\neq{\omega}}G^{{\omega}^{\prime}}_{{\omega}}\log G^{{\omega}^{\prime}}_{{\omega}}$ (A.85)
@@
Therefore, for any normalized state vector ${\hat{\bm{\varphi}}}\in{\mathcal{H}}$, we have
$\tilde{\bm{L}}{\widetilde{\mathbf{G}}}{\hat{\bm{\varphi}}}=\tilde{\bm{L}}{\widetilde{\mathbf{G}}}{\varphi}^{\omega}{\hat{\bm{e}}}_{{\omega}}\geq\sum_{\omega}{\varphi}^{\omega}\cdot(-K\log K)=-K\log K.$ (A.89)
@@
$E\left[\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L({\omega}_{t},{\omega}_{t}^{+})\Bigm{|}{\omega}_{0}={\omega}\right]$ $=KT+\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \tilde{\bm{L}}{\widetilde{\mathbf{G}}}e^{Gt}{\hat{\bm{e}}}_{{\omega}}$ (A.90)
@@ Appendix B Proofs in Section 3
${\hat{\bm{\varphi}}}(\infty):=\lim_{t\to\infty}{\hat{\bm{\varphi}}}(t)$ (B.1)
@@
The proof is not as innocent as its result looks to be. The proof consist of two parts: (1) we use the Gershgorin circle theorem to prove that all eigenvalues of ${{\mathbf{G}}}$ have negative real parts, or they are exactly 00; (2) we use Jordan normal form to show that $e^{{{\mathbf{G}}}t}$ converges as $t\to\infty$.
@@
Suppose our configuration set ${\Omega}$ has size $m$. Then, we can assign each ${\omega}\in{\Omega}$ a unique integer in $[m]:=\{1,2,\dots,m\}$. Thus, we find a matrix representation $G$ of our generator ${{\mathbf{G}}}$. We may as well use $G_{{\omega}}^{{\omega}^{\prime}}$ to denote the ${\omega}$-th row and ${\omega}^{\prime}$-th column of the matrix.
@@
As a result, the disc $D(G_{{\omega}}^{{\omega}},R_{j})$ must lie in the non-positive half of the complex plane, and the only point in the disc with non-negative real part is exactly 00.
@@
${\hat{\bm{\varphi}}}(t)=e^{{{\mathbf{G}}}t}{\hat{\bm{e}}}_{{\omega}}$ (B.14)
represents a probability distribution given any basis vector ${\hat{\bm{e}}}_{{\omega}}$. Therefore, $e^{Gt}$ is always bounded by a constant.
@@
Hence, we deduce that $e^{Jt}$ converges as $t\to\infty$, and thus the convergence of $e^{Gt}=Pe^{Jt}P^{-1}$. This implies the existence of
${\hat{\bm{\varphi}}}(\infty):=\lim_{t\to\infty}{\hat{\bm{\varphi}}}(t)=\lim_{t\to\infty}e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}(0).$ (B.15)
@@
Corollary 3.4 (${\hat{\bar{\bm{\varphi}}}}$ and ${\bar{\gamma}}$ Always Exist). Given that the configuration set ${\Omega}$ is finite, the averaged state ${\hat{\bar{\bm{\varphi}}}}$ and the objective value ${\bar{\gamma}}$ always exist. Moreover, we have
${\hat{\bar{\bm{\varphi}}}}={\hat{\bm{\varphi}}}(\infty)\qquad\text{and}\qquad{\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (B.16)
@@
Proposition 3.3 states that ${\hat{\bm{\varphi}}}(t)$ is converging in a Hilbert space ${\mathcal{H}}$. Thus, its mean would converge to the same limit, i.e., ${\hat{\bar{\bm{\varphi}}}}={\hat{\bm{\varphi}}}(\infty)$, and thus always exists.
@@
${\bar{\gamma}}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bm{\varphi}}}(t)={\bm{\Gamma}}{\widetilde{\mathbf{G}}}\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\hat{\bm{\varphi}}}(t)={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}},$ (B.17)
@@
Theorem 3.7 (Existence of Objective-Driven Dynamics: A Case Study). Consider the two-entity formulation described above for an ergodic system. If the objective operator ${\bm{\Gamma}}:{\widetilde{\mathcal{H}}}\to{\mathbb{R}}_{+}$ is non-negative and $\forall{{{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}}:\min_{{{\mathbf{M}}}\in{{\mathcal{P}}}_{x}}{\bar{\gamma}}({{\mathbf{M}}})=0$ for a finite ${{\mathcal{P}}}_{x}$, then there exists generator ${{\mathbf{M}}}^{\star}$ that simulates the minimization, such that
@@
:=×${\mathcal{M}}^{\star}:={\mathcal{M}}\times[m].$ (B.19)
However, when ${{\mathbf{M}}}^{\star}$ is simulating a ${{\mathbf{M}}}$, we may still work with the corresponding Hilbert space ${\widetilde{\mathcal{H}}}$. Then, we can construct the new generator ${{\mathbf{M}}}^{\star}$ as the following. For each ${\omega}\to{\omega}^{+}$, if its objective signal is ${\bm{\Gamma}}{\hat{\bm{e}}}_{{\omega}^{+}{\omega}}=0$, then ${{\mathbf{M}}}^{\star}$ keeps simulating the current ${{\mathbf{M}}}$. If its objective signal is positive, then ${{\mathbf{M}}}^{\star}$ has a positive transition rate to other strategies in ${{\mathcal{P}}}_{x}$ by setting the current strategy $i$ to a uniformly random label in $j\in[m]$.
Next, we prove that this ${{\mathbf{M}}}^{\star}$ leads to the convergence to a stationary distribution with zero objective value. We know that $\forall{{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}$, there are some ${{\mathbf{M}}}_{i}\in{{\mathcal{P}}}_{x}$ such that ${\bar{\gamma}}({{\mathbf{M}}}_{i})=0$. Let ${\hat{\bar{\bm{\varphi}}}}_{i}$ denote the corresponding stationary distribution, and by definition ${\bar{\gamma}}({{\mathbf{M}}}_{i})={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}_{i}=0$.
Let ${\Omega}_{i}\subset{\Omega}$ be the subset of configurations where ${\hat{\bar{\bm{\varphi}}}}_{i}$ has positive measure, i.e., ${\bar{\varphi}}_{i}^{\omega}>0$ if and only if ${\omega}\in{\Omega}_{i}$. We can see that, by definition, ${\Omega}_{i}$ must be closed, meaning that any configuration in ${\Omega}_{i}$ can only transit to another configuration in ${\Omega}_{i}$. Moreover, this transition can never incur positive objective signal ${\bm{\Gamma}}$, since ${\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}_{i}=0$. The important property we can see is that ${\Omega}_{i}$ is absorbing, i.e., once the system configuration transits into this set, it never leaves, and it never incurs positive objective value.
@@
Therefore, under ${{\mathbf{M}}}^{\star}$, for any configuration of the system outside such a ${\Omega}_{i}$, it must have a path with positive transition rate leading to one of such ${\Omega}_{i}$. Since such ${\Omega}_{i}$ are all absorbing, the system eventually must stay in one of those forever, and thus the corresponding stationary state ${\hat{\bar{\bm{\varphi}}}}_{i}$ incurs zero objective value.
@@
Fact 3.11 (Another Representation of ${\Delta\mathcal{H}}$).
${\Delta\mathcal{H}}={\mathbf{\Pi}}\ {\widetilde{\mathcal{H}}}=\left\{{\hat{\bm{\varphi}}}\in{\mathcal{H}}\mid\sum_{{\omega}}{\varphi}^{\omega}=0\right\}.$ (B.20)
@@
Let us denote $({\Delta\mathcal{H}})^{\prime}:=\left\{{\hat{\bm{\varphi}}}\in{\mathcal{H}}\mid\sum_{{\omega}}{\varphi}^{\omega}=0\right\}$. By definition of the projection operator ${\mathbf{\Pi}}$ (equation 3.11), it is easy to see that ${\mathbf{\Pi}}{\widetilde{\mathcal{H}}}\subseteq({\Delta\mathcal{H}})^{\prime}$.
For the other direction, suppose $\dim({\mathcal{H}})=n$, and we can see that $({\Delta\mathcal{H}})^{\prime}$ is a $(n-1)$-dimensional subspace of ${\mathcal{H}}$. Let us fix ${\omega}$ and choose the following basis vector for consideration:
$\{{\hat{\bm{e}}}_{{\omega}^{\prime}}-{\hat{\bm{e}}}_{{\omega}}\}_{{\omega}^{\prime}\in{\Omega}}.$ (B.21)
The above is a set of $n-1$ linearly independent vectors (ignoring the zero), and they all lie in the subspace $({\Delta\mathcal{H}})^{\prime}$ whose dimension is also $n-1$. Therefore, the above set of vectors span the entire $({\Delta\mathcal{H}})^{\prime}$, i.e.,
$({\Delta\mathcal{H}})^{\prime}=\texttt{span}\left(\{{\hat{\bm{e}}}_{{\omega}^{\prime}}-{\hat{\bm{e}}}_{{\omega}}\}_{{\omega}^{\prime}\in{\Omega}}\right).$ (B.22)
@@
$({\Delta\mathcal{H}})^{\prime}=\texttt{span}\left(\{{\hat{\bm{e}}}_{{\omega}^{\prime}}-{\hat{\bm{e}}}_{{\omega}}\}_{{\omega}^{\prime}\in{\Omega}}\right)={\mathbf{\Pi}}\ \texttt{span}\left(\{{\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}}\}_{{\omega}^{\prime}\in{\Omega}}\right)\subseteq{\mathbf{\Pi}}\ {\widetilde{\mathcal{H}}}.$ (B.23)
@@
${{\mathbf{G}}}={\hat{\bm{e}}}_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}{{\check{\bm{\eta}}}}^{{\omega}},\qquad{\widetilde{\mathbf{G}}}=G^{{\omega}^{\prime}}_{{\omega}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}.$ (B.25)
@@
${\mathbf{\Pi}}\ {\widetilde{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}$ $={\mathbf{\Pi}}\ G^{{\omega}^{\prime}}_{{\omega}}{\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}}=\sum_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}\cdot({\hat{\bm{e}}}_{{\omega}^{\prime}}-{\hat{\bm{e}}}_{{\omega}})$ (B.26)
$=\sum_{{\omega}^{\prime}:{\omega}^{\prime}\neq{\omega}}G^{{\omega}^{\prime}}_{{\omega}}\cdot({\hat{\bm{e}}}_{{\omega}^{\prime}}-{\hat{\bm{e}}}_{{\omega}}).$ (B.27)
@@
(B.27$)=G^{{\omega}^{\prime}}_{{\omega}}{\hat{\bm{e}}}_{{\omega}^{\prime}}={{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}.$ (B.28)
@@
Given that the system is ergodic, there is a unique stationary state ${\hat{\bar{\bm{\varphi}}}}$. In this case, it is known that the dynamics converge to the stationary state with an exponential rate. Formally, for any initial state ${\hat{\bm{\varphi}}}(0)$, there exist some constant $K,\kappa>0$ such that
$\left\|e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}(0)-{\hat{\bar{\bm{\varphi}}}}\right\|\leq Ke^{-\kappa t}.$ (B.29)
@@
Next, note that ${\Delta\mathcal{H}}$ is the span of vectors like ${\hat{\bm{e}}}_{{\omega}^{\prime}}-{\hat{\bm{e}}}_{{\omega}}$, and we can see that any $\Delta{\hat{\bm{\varphi}}}\in{\Delta\mathcal{H}}$ can be expressed by the difference of two states (normalized vectors), i.e.,
$\Delta{\hat{\bm{\varphi}}}=Z\cdot({\hat{\bm{\varphi}}}_{1}-{\hat{\bm{\varphi}}}_{2}),$ (B.30)
where $Z>0$ and ${\hat{\bm{\varphi}}}_{1},{\hat{\bm{\varphi}}}_{2}\in{\mathcal{H}}$ are normalized states representing probability distribution.

### v2

@@ Appendix A Proofs in Section 2
$p_{T}({\omega}^{\prime}\mid{\omega})=\langle{\omega}^{\prime}\rvert e^{{{\mathbf{G}}}T}\lvert{\omega}\rangle=\int{\mathcal{D}}{\omega}\ \exp\left\{-\sum_{x}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t,x},w_{t,x}^{+})\right\}.$ (A.70)
@@
It corresponds to an operator $\bm{\Gamma}$$\widetilde{\mathcal{H}}$$\to{\mathbb{R}}$ defined as
${\bm{\Gamma}}:=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}\gamma_{{\omega}^{\prime}{\omega}}\langle{\omega}^{\prime}{\omega}\rvert,\qquad\text{where }\ \forall{\omega}:\gamma_{{\omega}{\omega}}=0.$ (A.72)
@@
$\mathbb{E}\left[\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \gamma(t,{\omega},{\omega}^{+})\Bigm|{\omega}_{0}={\omega}\right]=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}\lvert{\omega}\rangle.$ (A.73)
@@
$=\left(\sum_{{\omega}^{\prime\prime}}\gamma_{{\omega}^{\prime\prime}{\omega}}\langle{\omega}^{\prime\prime}{\omega}\rvert\right)\left(\sum_{{\omega}^{\prime}}\lvert{\omega}^{\prime}{\omega}\rangle G^{{\omega}^{\prime}}_{{\omega}}\right)\cdot{\Delta t}+o({\Delta t})$ (A.76)
$={\bm{\Gamma}}{\widetilde{\mathbf{G}}}\lvert{\omega}\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.77)
@@
$=\sum_{n=1}^{N}\sum_{{\omega}^{\prime}}\langle{\omega}^{\prime}\rvert e^{(n-1){{\mathbf{G}}}{\Delta t}}\lvert{\omega}\rangle\cdot\left({\bm{\Gamma}}{\widetilde{\mathbf{G}}}\lvert{\omega}^{\prime}\rangle\cdot{\Delta t}+o({\Delta t})\right)$ (A.80)
$=\sum_{n=1}^{N}{\Delta t}\cdot\left({\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{(n-1){{\mathbf{G}}}{\Delta t}}\lvert{\omega}\rangle+o({\Delta t})/{\Delta t}\right)$ (A.81)
$=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}\lvert{\omega}\rangle.$ (as ${\Delta t}\to 0$ and $N\to\infty$)
@@
$\tilde{\bm{L}}{\widetilde{\mathbf{G}}}\lvert{\omega}\rangle=-K\log K+H_{\omega}.$ (A.83)
@@
For any $\lvert{\omega}\rangle\in{\mathcal{H}}$, we have
$\tilde{\bm{L}}{\widetilde{\mathbf{G}}}\lvert{\omega}\rangle$ $=-\sum_{{\omega}^{\prime}:{\omega}^{\prime}\neq{\omega}}G^{{\omega}^{\prime}}_{{\omega}}\log G^{{\omega}^{\prime}}_{{\omega}}$ (A.85)
@@
Therefore, for any normalized state vector $\lvert\varphi\rangle$$\in{\mathcal{H}}$, we have
$\tilde{\bm{L}}{\widetilde{\mathbf{G}}}{\lvert\varphi\rangle}=\sum_{\omega}\tilde{\bm{L}}{\widetilde{\mathbf{G}}}{\varphi}^{\omega}\lvert{\omega}\rangle\geq\sum_{\omega}{\varphi}^{\omega}\cdot(-K\log K)=-K\log K.$ (A.89)
@@
$E\left[\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L({\omega}_{t},{\omega}_{t}^{+})\Bigm|{\omega}_{0}={\omega}\right]$ $=KT+\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \tilde{\bm{L}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}\lvert{\omega}\rangle$ (A.90)
@@ Appendix B Proofs in Section 3
$\lvert\varphi(\infty)\rangle:=\lim_{t\to\infty}\lvert\varphi(t)\rangle$ (B.1)
@@
The proof is not as innocent as its result looks to be. The proof consists of two parts: (1) we use the Gershgorin circle theorem to prove that all eigenvalues of $\mathbf{G}$}} have negative real parts, or they are exactly 0; (2) we use Jordan normal form to show that $e^{{{\mathbf{G}}}t}$ converges as $t\to\infty$.
@@
[…]
@@
As a result, the disc $D(G_{{\omega}}^{{\omega}},R_{j})$ must lie in the non-positive half of the complex plane, and the only point in the disc with non-negative real part is exactly 0.
@@
$\lvert\varphi(t)\rangle=e^{{{\mathbf{G}}}t}\lvert{\omega}\rangle$ (B.14)
represents a probability distribution given any basis vector $\lvert{\omega}\rangle$. Therefore, $e^{{{\mathbf{G}}}t}$ is always bounded by a constant.
@@
Hence, we deduce that $e^{Jt}$ converges as $t\to\infty$, and thus the convergence of $e^{{{\mathbf{G}}}t}=Pe^{Jt}P^{-1}$. This implies the existence of
$\lvert\varphi(\infty)\rangle:=\lim_{t\to\infty}\lvert\varphi(t)\rangle=\lim_{t\to\infty}e^{{{\mathbf{G}}}t}\lvert\varphi(0)\rangle.$ (B.15)
@@
Corollary 3.4 ($\lvert\bar{\varphi}\rangle$} and $\bar{\gamma}$} Always Exist). Given that the configuration set $\Omega$} is finite, the averaged state $\lvert\bar{\varphi}\rangle$} and the objective value $\bar{\gamma}$} always exist. Moreover, we have
${\lvert\bar{\varphi}\rangle}=\lvert\varphi(\infty)\rangle\qquad\text{and}\qquad{\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (B.16)
@@
Proposition 3.3 states that $\lvert\varphi(t)\rangle$ is converging in a Hilbert space $\mathcal{H}$}. Thus, its mean would converge to the same limit, i.e., ${\lvert\bar{\varphi}\rangle}=\lvert\varphi(\infty)\rangle$, and thus always exists.
@@
${\bar{\gamma}}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}\lvert\varphi(t)\rangle={\bm{\Gamma}}{\widetilde{\mathbf{G}}}\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \lvert\varphi(t)\rangle={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle},$ (B.17)
@@
Theorem 3.7 (Existence of Objective-Driven Dynamics: A Case Study). Consider the two-entity formulation described above for an ergodic system. If the objective operator $\bm{\Gamma}$$\widetilde{\mathcal{H}}$$\to{\mathbb{R}}_{+}$ is non-negative and $\forall{{{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}}:\min_{{{\mathbf{M}}}\in{{\mathcal{P}}}_{x}}{\bar{\gamma}}({{\mathbf{M}}})=0$ for a finite ${{\mathcal{P}}}_{x}$, then there exists generator $\mathbf{M}$$^{\star}$ that simulates the minimization, such that
@@
${\mathcal{M}}^{\star}:={\mathcal{M}}\times[m].$ (B.19)
However, when $\mathbf{M}$$^{\star}$ is simulating a $\mathbf{M}$}}, we may still work with the corresponding Hilbert space $\widetilde{\mathcal{H}}$}. Then, we can construct the new generator $\mathbf{M}$$^{\star}$ as the following. For each ${\omega}\to{\omega}^{+}$, if its objective signal is ${\bm{\Gamma}}\lvert{\omega}^{+}{\omega}\rangle=0$, then $\mathbf{M}$$^{\star}$ keeps simulating the current $\mathbf{M}$}}. If its objective signal is positive, then $\mathbf{M}$$^{\star}$ has a positive transition rate to other strategies in ${{\mathcal{P}}}_{x}$ by setting the current strategy $i$ to a uniformly random label in $j\in[m]$.
Next, we prove that this $\mathbf{M}$$^{\star}$ leads to the convergence to a stationary distribution with zero objective value. We know that $\forall{{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}$, there are some ${{\mathbf{M}}}_{i}\in{{\mathcal{P}}}_{x}$ such that ${\bar{\gamma}}({{\mathbf{M}}}_{i})=0$. Let ${\lvert\bar{\varphi}\rangle}_{i}$ denote the corresponding stationary distribution, and by definition ${\bar{\gamma}}({{\mathbf{M}}}_{i})={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}_{i}=0$.
Let ${\Omega}_{i}\subset{\Omega}$ be the subset of configurations where ${\lvert\bar{\varphi}\rangle}_{i}$ has positive measure, i.e., ${\bar{\varphi}}_{i}^{\omega}>0$ if and only if ${\omega}\in{\Omega}_{i}$. We can see that, by definition, ${\Omega}_{i}$ must be closed, meaning that any configuration in ${\Omega}_{i}$ can only transit to another configuration in ${\Omega}_{i}$. Moreover, this transition can never incur positive objective signal $\bm{\Gamma}$}, since ${\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}_{i}=0$. The important property we can see is that ${\Omega}_{i}$ is absorbing, i.e., once the system configuration transits into this set, it never leaves, and it never incurs positive objective value.
@@
Therefore, under $\mathbf{M}$$^{\star}$, for any configuration of the system outside such a ${\Omega}_{i}$, it must have a path with positive transition rate leading to one of such ${\Omega}_{i}$. Since such ${\Omega}_{i}$ are all absorbing, the system eventually must stay in one of those forever, and thus the corresponding stationary state ${\lvert\bar{\varphi}\rangle}_{i}$ incurs zero objective value.
@@
Fact 3.11 (Another Representation of $\Delta\mathcal{H}$}). Write ${\lvert\varphi\rangle}=\sum_{{\omega}}{\varphi}^{\omega}\lvert{\omega}\rangle$, and we have
${\Delta\mathcal{H}}={\mathbf{\Pi}}\ {\widetilde{\mathcal{H}}}=\left\{{\lvert\varphi\rangle}\in{\mathcal{H}}\mid\sum_{{\omega}}{\varphi}^{\omega}=0\right\}.$ (B.20)
@@
Let us denote $({\Delta\mathcal{H}})^{\prime}:=\left\{{\lvert\varphi\rangle}\in{\mathcal{H}}\mid\sum_{{\omega}}{\varphi}^{\omega}=0\right\}$. By definition of the projection operator $\mathbf{\Pi}$} (equation 3.11), it is easy to see that ${\mathbf{\Pi}}{\widetilde{\mathcal{H}}}\subseteq({\Delta\mathcal{H}})^{\prime}$.
For the other direction, suppose $\dim({\mathcal{H}})=n$, and we can see that $({\Delta\mathcal{H}})^{\prime}$ is a (n1)(n-1)-dimensional subspace of $\mathcal{H}$}. Let us fix $\omega$} and choose the following basis vector for consideration:
$\{\lvert{\omega}^{\prime}\rangle-\lvert{\omega}\rangle\}_{{\omega}^{\prime}\in{\Omega}}.$ (B.21)
The above is a set of n1n-1 linearly independent vectors (ignoring the zero), and they all lie in the subspace $({\Delta\mathcal{H}})^{\prime}$ whose dimension is also n1n-1. Therefore, the above set of vectors span the entire $({\Delta\mathcal{H}})^{\prime}$, i.e.,
$({\Delta\mathcal{H}})^{\prime}=\texttt{span}\left(\{\lvert{\omega}^{\prime}\rangle-\lvert{\omega}\rangle\}_{{\omega}^{\prime}\in{\Omega}}\right).$ (B.22)
@@
$({\Delta\mathcal{H}})^{\prime}=\texttt{span}\left(\{\lvert{\omega}^{\prime}\rangle-\lvert{\omega}\rangle\}_{{\omega}^{\prime}\in{\Omega}}\right)={\mathbf{\Pi}}\ \texttt{span}\left(\{\lvert{\omega}^{\prime}{\omega}\rangle\}_{{\omega}^{\prime}\in{\Omega}}\right)\subseteq{\mathbf{\Pi}}\ {\widetilde{\mathcal{H}}}$. (B.23)
@@
${{\mathbf{G}}}=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}\lvert{\omega}^{\prime}\rangle G^{{\omega}^{\prime}}_{{\omega}}\langle{\omega}\rvert,\qquad{\widetilde{\mathbf{G}}}=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}G^{{\omega}^{\prime}}_{{\omega}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}.$ (B.25)
@@
${\mathbf{\Pi}}\ {\widetilde{\mathbf{G}}}\lvert{\omega}\rangle$ $=\sum_{{\omega}^{\prime}}{\mathbf{\Pi}}\ G^{{\omega}^{\prime}}_{{\omega}}\lvert{\omega}^{\prime}{\omega}\rangle=\sum_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}\cdot(\lvert{\omega}^{\prime}\rangle-\lvert{\omega}\rangle)$ (B.26)
$=\sum_{{\omega}^{\prime}:{\omega}^{\prime}\neq{\omega}}G^{{\omega}^{\prime}}_{{\omega}}\cdot(\lvert{\omega}^{\prime}\rangle-\lvert{\omega}\rangle).$ (B.27)
@@
[…]
@@
Given that the system is ergodic, there is a unique stationary state $\lvert\bar{\varphi}\rangle$}. In this case, it is known that the dynamics converge to the stationary state with an exponential rate. Formally, for any initial state $\lvert\varphi(0)\rangle$, there exist some constant $K,\kappa>0$ such that
$\left\|e^{{{\mathbf{G}}}t}\lvert\varphi(0)\rangle-{\lvert\bar{\varphi}\rangle}\right\|\leq Ke^{-\kappa t}.$ (B.29)
@@
Next, note that $\Delta\mathcal{H}$} is the span of vectors like $\lvert{\omega}^{\prime}\rangle-\lvert{\omega}\rangle$, and we can see that any $\Delta{\lvert\varphi\rangle}\in{\Delta\mathcal{H}}$ can be expressed by the difference of two states (normalized vectors), i.e.,
$\Delta{\lvert\varphi\rangle}=Z\cdot({\lvert\varphi\rangle}_{1}-{\lvert\varphi\rangle}_{2}),$ (B.30)
where $Z>0$ and ${\lvert\varphi\rangle}_{1},{\lvert\varphi\rangle}_{2}\in{\mathcal{H}}$ are normalized states representing probability distribution.